OpenStudy (anonymous):

f(x)=e^sqrt(3x) differentiate

OpenStudy (anonymous):

you have to use chain rule \[y=e^{\sqrt{3x}}\\ y'=e^\sqrt{3x}{d\over dx}(\sqrt{3x}) \] and continue

OpenStudy (anonymous):

is this the answer 3/2sqrtx (e^3x

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you still need \(\sqrt{3x}\) in the exponent

OpenStudy (anonymous):

you also need the derivative of \(\sqrt{3x}=\frac{3}{2\sqrt{3x}}\)